OpenStudy (anonymous):

Suppose that the amount of algae in a pond doubles every 4 hours. If the pond initially contains 90 pounds of algae, how much algae will be in the pond after 12 hours

OpenStudy (zpupster):

53 zpupster 0P = Noe^(kt) where P is amount after doubling No is the initial amount . No is 90 pounds P is 180 pounds t = 4 hrs . 180= 90e^(4k) 2= e^(4k) ln(2) = 4k ln(2)/4 = k 0.173287 =k F(12) = 1000e(0.173287*12)